# Quadrature Over-the-Air-Computing for Multimodal Dual-Stream Signal Processing

### v1

@@ IV Performance Analysis
It should be noted that a relative performance variation among the different function combinations can be seen, but this is not a remnant of the combiner design, but reflects the underlying numerical behavior of the pre-/post-processing mappings of the nomographic functions. For example, the product function involves logarithmic and exponential operations, which are inherently more sensitive to noise than the identity mappings used for the summation and mean. Nevertheless, the results validate the ability of the proposed Q-OTAC for multimodal function aggregation, and also motivates a future consideration to a joint optimization of the precoders and the combiners to mitigate this effect under the assumption of a feedback loop between the EDs and the AP.
Figure 3: The NMSE CDF for different multimodal dual-stream Q-OTAC scenarios (simultaneous sum and product, and simultaneous arithmetic mean and weighted average) for a system with $N=20,K=20$ and $\mathrm{SNR}=15\mathrm{dB}$.
@@
Figure 4: Summation function performance comparison of the proposed dual-stream Q-OTAC against the conventional single-stream OTAC, with $N=20$ and $K=20$.
[…]
V Conclusion and Future Works
We proposed Q-OTAC, the first over-the-air computing framework to enable dual-function aggregation in a single transmission via a novel IQ-decoupled MMSE combiner. Simulation results confirm its effectiveness in both heterogeneous function settings, achieving twice the computation rate and improved average accuracy over the single-stream OTAC, supporting multimodal and scalable B5G OTAC scenarios. Future works will investigate extended designs of the proposed combiner, also jointly with the precoder optimization, consider the system in under-/over-loaded scenarios, and most importantly incoporate the Q-OTAC with ICC and ISCC frameworks.

### v2

@@ 4 Performance Analysis
A relative performance variation among the different function combinations can also be observed, but this is not a remnant of the combiner design, but reflects the underlying numerical behavior of the pre-/post-processing of the nomographic functions. For example, the product function involves logarithmic and exponential operations, which are inherently more sensitive to noise than the identity mappings used for the summation and mean.
@@
Fig. 3: Summation performance of the proposed Q-OTAC against the conventional OTAC, with $N=20$, $K=20$.
Next, Figure 3 compares the performance of the proposed dual-stream Q-OTAC with the conventional single-stream OTAC scheme, evaluated in terms of the total NMSE per computation via the CDF and average NMSE over SNR, in the respective subfigures. The results demonstrate that Q-OTAC achieves a noticable improved total NMSE per computation, over the entire range of both the CDF and SNR plots with an approximate gain of $\penalty 10000\ 5\text{dB}$ thanks to its novel and efficient utilization of both the IQ components of the complex signal to perform computing, as opposed to only a single real component in the SotA, thereby doubling the signal space and effective computation rate.
5 Conclusion and Future Works
We proposed the quadrature over-the-air computing (Q-OTAC), the first OTAC framework to enable dual-function aggregation in a single transmission via a novel IQ-decoupled MMSE combiner. Simulation results confirm its effectiveness in heterogeneous function settings, achieving twice the computation rate and improved average accuracy over the single-stream OTAC, supporting multimodal and scalable B5G OTAC scenarios. Future works will investigate extended combiner designs, jointly with the precoder optimization considerations, and incoporate ICC/ISCC approaches.